OpenStudy (anonymous):

4 over (3x)^-2 4 over (3x)^-2 @Mathematics

OpenStudy (saifoo.khan):

\[\Large \frac{4}{(3x)^{-2}} \to4 \times 9x^2 \to 36x^2\]

OpenStudy (alfie):

\[\frac{4}{{{{(3x)}^{ - 2}}}} \to 4*\frac{1}{{{{(3x)}^{ - 2}}}} \to 4*9{x^2} \to 36{x^2}\]